How do you find profit maximizing price given a marginal cost curve and a demand curve?
Answer #1

From the demand curve, we find MR and equate it with MC. We get equilibrium quantity, when we put value of Q in demand curve equation we get equilibrium price.

Example; Demand function, Q = 40 - P and MC = 10

Q = 40 - P

P = 40 - Q

TR = PQ = (40 - Q)Q = 40Q - Q2

MR = 40 - 2Q

MC = 10

Equilibrium condition; MR = MC

40 - 2Q = 10

2Q = 30

Q = 15 units

P = 40 - 15

P = $ 25